# Spatial-Aware Conformal Prediction for Trustworthy Hyperspectral Image Classification

### v1

@@ III Validity of CP on HSI
Exchangeability and validity of CP on HSI classification
As outlined in Theorem 1, CP necessitates that calibration and test samples be exchangeable. However, HSI classifiers typically operate on pixel cuboids as the fundamental unit of training, rather than individual pixels, and thus require access to the entire HSI during training. Therefore, when partitioning $\mathcal{D}_{\mathcal{I}}$ into $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{cal}}$, and $\mathcal{D}_{\text{test}}$, cuboids in $\mathcal{D}_{\text{train}}$ may overlap with those in $\mathcal{D}_{\text{cal}}$ or $\mathcal{D}_{\text{test}}$ due to shared pixels, which potential breaks the validity of CP. In this section, we demonstrate that, under a general permutation-invariance condition, the exchangeability of the non-conformity scores remains valid even when both calibration and test data are used during training. This theoretical analysis conclusively establishes the validity of applying CP to HSI classification.
@@
The proof of Theorem 2 is detailed in Appendix A. This theorem demonstrates that exchangeability is preserved and establishes the validity of applying CP to HSI classification. We also empirically demonstrate the coverage guarantee in our experiments, as detailed in Table II. A similar analysis is also discussed in prior work [30], which establishes the applicability of CP to Graph Neural Networks. Building on this theoretical foundation, we propose a framework to equip HSI classifiers with reliable conformal prediction sets.
Conformal HSI Classification
For an HSI $\mathcal{I}$ with a random subset of labels known, the following steps outline how to generate prediction sets for the remaining unlabeled pixels to include their true labels with a coverage rate of $\alpha$:
@@
    Define the score function $S$ based on the classifier $\hat{\pi}_{\theta}$ and compute the non-conformity scores $s_{i}=S(B_{i},y_{i})$ for each sample in the calibration set $\mathcal{D}_{\text{cal}}$.
@@
Theorem 2 implies that the non-conformity scores of $\mathcal{D}_{\text{cal}}$ and $\mathcal{D}_{\text{test}}$ are exchangeable. Consequently, it can be inferred that the prediction set $\mathcal{C}_{1-\alpha}(B_{i};\tau)$ will include the true label of the central pixel of $B_{i}$ with a probability of $1-\alpha$. For the non-conformity score function $S$, we can employ APS, RAPS or SAPS. However, these score functions depend exclusively on softmax probabilities and label ranks, neglecting the crucial spatial information inherent in HSIs. This observation motivates the exploration of integrating spatial information into the score calculation to enhance CP performance.
@@ IV-A Motivation
Numerous studies have demonstrated that leveraging spatial features can significantly enhance HSI classification accuracy [35, 36]. However, in the context of CP, existing non-conformity scores are computed solely based on the output conditional probability of the central pixel, which results in a considerable neglect of spatial information. Below, we present an analysis that illustrates the impact of incorporating spatial information into the calculation of non-conformity scores.
Given a non-conformity score function $S$ based on an HSI classifier, the score for a pair $(B_{i},y_{i})$ is $S(B_{i},y_{i})$. To incorporate information from neighboring cuboids, we define the adjusted score function $\tilde{S}(B_{i},y_{i})$ as follows:
@@
where $\mathcal{N}_{i}$ denotes the set of neighboring cuboids around the central pixel of $B_{i}$. Intuitively speaking, since neighboring pixels in HSI often belong to the same class, if $y_{i}$ is indeed the true label for $B_{i}$, it is likely that $y_{i}$ is also the true label for $B_{j}\in\mathcal{N}_{i}$. Consequently, the term $\frac{1}{|\mathcal{N}_{i}|}\sum_{B_{j}\in\mathcal{N}_{i}}S(B_{j},y_{i})$ will generally be small. Conversely, if $y_{i}$ is not the correct label for $B_{i}$, this summation is expected to be larger. Thus, the adjusted score $\tilde{S}(B_{i},y_{i})$ will increase less for correct instance-label pairs and more for incorrect ones. The advantages of this adjustment are further elucidated by the following proposition.
@@
The proof of Proposition 1 is provided in Appendix B. It suggests that the probability of a randomly selected score from $\mathcal{D}_{\text{cal}}$ exceeding a randomly selected one from $\hat{\mathcal{D}}_{\text{test}}$ serves as an indicator of CP performance: a lower probability indicates more efficient prediction sets. For the adjusted score $\tilde{S}$ in Eq. (14), which incorporates aggregation within neighborhoods, there is typically a larger increase in scores for $\hat{\mathcal{D}}_{\text{test}}$ compared to those in $\mathcal{D}_{\text{cal}}$. This discrepancy arises because the instance-label pairs in $\mathcal{D}_{\text{cal}}$ are always correctly matched, whereas $\hat{\mathcal{D}}_{\text{test}}$ predominantly contains mismatched ones. Thus, compared to standard CP, the incorporation of spatial information (via Eq. (14)) is expected to reduce the probability defined in Proposition 1, resulting in more efficient prediction sets.
@@ IV-B Methodology
1:  Input: Training set $\mathcal{D}_{\text{train}}$, calibration set $\mathcal{D}_{\text{cal}}$, test cuboid $B_{n+1}$, error rate $\alpha$, and parameters $k$ and $\lambda$.
2:  Output: Conformal prediction set $\mathcal{C}_{1-\alpha}(B_{n+1};\hat{\tau})$.
3:  Step 1: Train a deep HSI classifier $\hat{\pi}_{\theta}$ using $\mathcal{D}_{\text{train}}$.
4:  Step 2: Define the score function $S$ based on $\hat{\pi}_{\theta}$.
5:  Step 3: Compute the scores for $\mathcal{D}_{\text{cal}}$ and aggregate the results using the Score Aggregation Operator defined in Eq. (15). Denote the aggregated scores as $\{\hat{s}_{i}\}_{i=1}^{{\scriptstyle|\mathcal{D}_{\text{cal}}|}}$.
6:  Step 4: Determine the threshold $\hat{\tau}$ by:
@@
7:  Step 5: For $B_{n+1}$, the prediction set is derived as:
@@
To demonstrate that SACP can produce trustworthy prediction sets, we also theoretically establish that SACP ensures coverage, which is formalized in the following proposition:
@@
  • Model-agnostic: SACP is a post-hoc approach that can be integrated into any pre-trained HSI classifier without requiring modifications to the model itself.
@@
  • Computationally efficient: SACP maintains computational efficiency comparable to that of standard CP (see the discussion in Section V-C).
@@ V-A Experimental Setup
We utilize the commonly used benchmarks Indian Pines (IP), Pavia University (PU), and Salinas (SA) datasets for our evaluations [^2] These datasets are available at https://www.ehu.eus/ccwintco/index.php/Hyperspectral_Remote_Sensing_Scenes. Specifically, IP consists of $145\times 145$ pixels and 200 valid spectral bands, representing 16 different land-cover classes. PU includes $610\times 340$ pixels with 103 usable spectral bands, covering 9 land-cover categories. SA comprises $512\times 217$ pixels with 204 effective spectral bands, corresponding to 16 ground-truth classes. For all these HSI datasets, we first perform random sampling of HSI cuboids to create the training set. The remaining data is then arbitrarily divided into calibration and test sets.
@@
Model $\mathcal{D}_{\mathcal{I}}$ $|\mathcal{D}_{\text{train}}|$ $|\mathcal{D}_{\text{cal}}|$ $|\mathcal{D}_{\text{test}}|$
1D-CNN IP 250 4996 5003
PU 103 21336 21337
SA 244 26943 26942
3D-CNN IP 128 5057 5064
PU 52 21362 21362
SA 122 27003 27004
HybridSN IP 128 5057 5064
PU 52 21362 21362
SA 122 27003 27004
SSTN IP 128 5057 5064
PU 52 21362 21362
SA 122 27003 27004
Models
To validate the generalizability of SACP, we employ four widely-used HSI classifiers, incorporating both CNN-based and Transformer-based models with various scales and architectures. For CNN-based models, we select 1D-CNN [37], 3D-CNN [38], and HybridSN [39]. The 1D-CNN and 3D-CNN differ in the dimensions of their convolutional kernels, while HybridSN utilizes multi-dimensional kernels to enhance feature extraction. For the Transformer-based model, we use SSTN [40]. All HSI classifiers are initially pre-trained on the three datasets and then evaluated on the corresponding test data. Table I details the sizes of the training, calibration, and test datasets for each dataset across different models. Due to its relatively poor performance, the 1D-CNN is allocated a larger amount of training data compared to the other three models. For clarity, Table II demonstrates the Overall Accuracy (OA) and Average Accuracy (AA) results for the pre-trained HSI classifiers across the three datasets. The definitions of OA and AA are given as follows:
where $\mathcal{D}_{\text{test}}:=\{(B_{i},y_{i})\}_{i=1}^{|\mathcal{D}_{\text{test}}|}$, $\hat{y}_{i}=\mathrm{argmax}(\hat{\pi}_{\theta}(B_{i}))$, $\mathds{1}$ is the indicator function, and $j\in\{1,\dots,|\mathcal{Y}|\}$. Note: In HSI classifier pre-training, the primary focus is not on maximizing their performance metrics such as OA and AA, but rather on achieving model diversity. Our goal is to validate the effectiveness—i.e., the generalizability—of our approach across different model architectures and training processes, irrespective of their performance levels. Given that our primary goal is not to optimize classification performance, we utilize only a minimal amount of data for training these HSI classifiers.
TABLE II: The results of OA and AA on various datasets of different classifiers.
@@
TABLE III: Performance comparison of prediction sets generated by SACP and standard CP across various HSI classifiers and datasets with different error rates $\alpha$. Bold numbers indicate that SACP is superior. $\downarrow$ indicates that smaller values are better.
@@
V-B Main Results
SACP yields efficient prediction sets
We compare the performance of prediction sets generated by standard CP and SACP across various HSI classifiers and datasets with different error rates $\alpha$. The results, presented in Table III, demonstrate that SACP consistently maintains the desired coverage of $1-\alpha$, empirically validating its coverage guarantee as outlined in Proposition 2. In other words, for each pixel, the true labels are reliably included in their prediction set with a user-specified probability. Formally, this is expressed as:
$\mathbb{P}(Y_{\text{test}}\in\mathcal{C}(X_{\text{test}})))\geq 1-\alpha.$ (18)
Additionally, SACP consistently produces more efficient prediction sets while maintaining a satisfactory conditional coverage rate. For instance, when applying 3D-CNN to the IP dataset, SACP reduces the Size from 5.73 for APS to 3.27, representing a 42.9% decrease. Similarly, for RAPS on the same dataset with 1D-CNN, SACP lowers the Size from 3.64 to 2.26, reflecting a 37.9% reduction. These improvements are observed across different error rates, non-conformity score functions, model architectures, and datasets, highlighting the robustness and generalizability of SACP. It is important to note that, unlike traditional classification tasks, SAPS and RAPS do not consistently outperform APS in the context of HSI classification. This further underscores the challenges associated with CP in HSI classification. Nevertheless, SACP consistently enhances the performance of various non-conformity score functions, affirming its effectiveness.
SACP consistently reduces prediction set size across various pixels
To further demonstrate the effectiveness of SACP, we visualize the size of prediction sets for each pixel in $\mathcal{D}_{\text{cal}}$ and $\mathcal{D}_{\text{test}}$ in Figure 2. This experiment is conducted on the IP and PU datasets using SSTN, with APS as the base score function and $\alpha$ set to 0.05. In the visualization, brighter pixels represent larger prediction sets. The results clearly indicate that SACP significantly reduces the size of prediction sets compared to standard CP. Moreover, standard CP often produces unusually large prediction sets, even for interior pixels, as evidenced in the lower-left corner of Figure 2(d). Typically, edge pixels, which are more challenging to predict, tend to have larger prediction sets, while interior pixels have smaller sets. SACP effectively addresses this issue by consistently reducing prediction set sizes for both edge and interior pixels. This improvement underscores the superior performance of our approach in conformal HSI classification.
Figure 2: Visualizations of the average size of prediction sets for each pixel in $\mathcal{D}_{\text{cal}}$ and $\mathcal{D}_{\text{test}}$ are presented for the (a) IP and (e) PU datasets. The true labels for each class are shown in (b) and (f). The brightness of each pixel indicates the size of its prediction set, with brighter pixels representing larger sets. Panels (c) and (g) display the set sizes with SACP, while panels (d) and (h) show the set sizes with standard CP.
SACP is insensitive to hyperparameters
In SACP, two hyperparameters are crucial: $\lambda$ and $k$. The parameter $\lambda$ serves as a trade-off, balancing the influence of the neighborhood, while $k$ specifies the number of iterations. Increasing $k$ allows SACP to incorporate spatial information from progressively larger neighborhoods. This section examines the impact of variations in $\lambda$ and $k$ on the performance of SACP. The experiment is conducted on all selected datasets using SSTN, with APS as the score function and $\alpha$ set to 0.05. The results are illustrated in Figure 3. Figure 3a demonstrates that SACP consistently outperforms the baseline across different values of $\lambda$, with the size of prediction sets decreasing as $\lambda$ increases. Figure 3b illustrates that SACP consistently achieves superior performance as the number of iterations $k$ changes. Overall, the results suggest that SACP is relatively insensitive to variations in $\lambda$ and $k$, consistently surpassing the baseline. For simplicity, we employ a moderate range of hyperparameters, consistently setting $\lambda=0.5$ and $k=1$ throughout the experiments.
@@ V-B Main Results
V-C Discussion
SACP is robust to the size of calibration set
As shown in Eq. (4), the calibration set is used to determine the threshold $\tau$ corresponding to the error rate of $1-\alpha$. A precise estimate of $\tau$ is essential. In this work, all experiments are conducted with a fixed ratio for the calibration set size, specifically $|\mathcal{D}_{\text{cal}}|\approx 0.5\times|\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}|$. In this section, we investigate whether SACP is sensitive to the size of the calibration set $|\mathcal{D}_{\text{cal}}|$. Specifically, we exam whether Coverage and Size remain satisfactory when varying the ratio of the calibration set size. The experiment is conducted on all three datasets using SSTN, with APS as the base score function and the error rate $\alpha$ set to $0.05$. The ratio of calibration samples is defined as $\gamma=(|\mathcal{D}_{\text{cal}}|/|\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}|)$. $\gamma$ ranges from 0.1 to 0.6 in increments of 0.05. The experimental results are presented in Table IV. Coverage consistently maintains $1-\alpha$ as the calibration set size decreases relative to $|\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}|$, even for very small values of $\gamma$. Additionally, the size of the prediction sets remains stable, only marginally influenced by changes in the calibration set size, and fluctuates within a narrow range. These findings collectively demonstrate that SACP exhibits significant robustness to variations in the calibration set size.

### v2

@@ II-B HSI Classification
III Validity of Conformal Prediction on HSI
III-A Theoretical Analysis
As outlined in Theorem 1, conformal prediction requires that calibration and test samples be exchangeable. However, HSI classifiers typically operate on pixel cuboids as the fundamental unit of training, rather than individual pixels, and thus require access to the entire HSI during training. Therefore, when partitioning $\mathcal{D}_{\mathcal{I}}$ into $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{cal}}$, and $\mathcal{D}_{\text{test}}$, cuboids in $\mathcal{D}_{\text{train}}$ may overlap with those in $\mathcal{D}_{\text{cal}}$ or $\mathcal{D}_{\text{test}}$ due to shared pixels, which potential breaks the validity of conformal prediction. In this section, we demonstrate that, under a general permutation-invariance condition, the exchangeability of the non-conformity scores remains valid even when pixels from both the calibration and test sets are used during training. This theoretical analysis conclusively establishes the validity of applying conformal prediction to HSI classification.
@@ III-A Theoretical Analysis
The proof of Theorem 2 is provided in Appendix A. This theorem establishes that exchangeability is preserved, thereby validating the application of conformal prediction to HSI classification. Specifically, it guarantees that the generated prediction sets will include the true label with a user-specified coverage probability. Additionally, we empirically demonstrate this coverage guarantee in our experiments, as detailed in Table II. Building on this theoretical foundation, we propose a conformal procedure that equips any trained HSI classifier with reliable prediction sets, ensuring that these sets contain the true labels with a user-specified probability.
III-B Conformal HSI Classification
For an HSI $\mathcal{I}$ with a randomly selected subset of known labels, the following steps outline the process for generating conformal prediction sets with a coverage rate of $1-\alpha$:
@@ III-B Conformal HSI Classification
    Define the score function $S$ based on the HSI classifier $\hat{\pi}_{\theta}$ and compute the non-conformity scores $s_{i}=S(B_{i},y_{i})$ for each sample in the calibration set $\mathcal{D}_{\text{cal}}$.
@@
Theorem 2 indicates that the non-conformity scores of $\mathcal{D}_{\text{cal}}$ and $\mathcal{D}_{\text{test}}$ are exchangeable. Thus, according to Theorem 1, it follows that the prediction set $\mathcal{C}_{1-\alpha}(B_{i};\tau)$ will include the true label of the central pixel of $B_{i}$ with a probability of $1-\alpha$. For the non-conformity score function $S$, we can utilize any well-defined non-conformity score functions, such as APS, RAPS, or SAPS. However, these score functions rely solely on softmax probabilities and label ranks, overlooking the critical spatial information inherent in HSIs. This observation motivates the exploration of integrating spatial information into the score calculation for enhanced conformal prediction.
@@ IV-A Motivation
Numerous studies have demonstrated that leveraging spatial information can significantly enhance HSI classification accuracy [35, 36]. However, in the context of conformal prediction, existing non-conformity scores are computed solely based on the output conditional probability of the central pixel and label rankings, resulting in a considerable neglect of spatial information. In the following, we present an theoretical analysis illustrating the impact of incorporating spatial information into the calculation of non-conformity scores.
Given a non-conformity score function $S$ based on an HSI classifier $\hat{\pi}_{\theta}$, the score for instance-label pair $(B_{i},y_{i})$ is $S(B_{i},y_{i})$. To incorporate information from neighboring cuboids, we define the refined score function $\tilde{S}(B_{i},y_{i})$ by:
@@
where $\mathcal{N}_{i}$ denotes the set of neighboring cuboids around the central pixel of $B_{i}$. Intuitively speaking, since neighboring pixels in a given HSI often belong to the same class, if $y_{i}$ is the true label for $B_{i}$, it is likely that $y_{i}$ is also the true label for $B_{j}\in\mathcal{N}_{i}$. Consequently, the term $\frac{1}{|\mathcal{N}_{i}|}\sum_{B_{j}\in\mathcal{N}_{i}}S(B_{j},y_{i})$ will generally be small. Conversely, if $y_{i}$ is not the correct label for $B_{i}$, this summation is expected to be larger. Therefore, compared to $S(B_{i},y_{i})$, the refined score $\tilde{S}(B_{i},y_{i})$ increases modestly for correct instance-label pairs and substantially for incorrect ones. To clarify, the advantages of this aggregation are formally demonstrated in the following proposition.
@@
The proof of Proposition 1 is provided in Appendix B. This proposition suggests that the probability of a randomly selected score from $\mathcal{D}_{\text{cal}}$ exceeding a randomly selected one from $\hat{\mathcal{D}}_{\text{test}}$ serves as an indicator of the efficiency of conformal predictors: a lower probability indicates more efficient prediction sets. For the refined score $\tilde{S}$ in Eq. (4), which incorporates aggregation within neighborhoods, scores in $\hat{\mathcal{D}}_{\text{test}}$ typically show a larger increase than those in $\mathcal{D}_{\text{cal}}$. This discrepancy occurs because the instance-label pairs in $\mathcal{D}_{\text{cal}}$ are always correctly matched, while the defined set $\hat{\mathcal{D}}_{\text{test}}$ mostly contains mismatched pairs. Therefore, compared to standard conformal prediction, the integration of spatial information (via Eq. (4)) is expected to lower the probability defined in Proposition 1, resulting in more efficient prediction sets.
@@ IV-B Methodology
  Input: Training set $\mathcal{D}_{\text{train}}$, calibration set $\mathcal{D}_{\text{cal}}$, test cuboid $B_{n+1}$, error rate $\alpha$, and parameters $k$ and $\lambda$.
  Output: Conformal prediction set $\mathcal{C}_{1-\alpha}(B_{n+1};\hat{\tau})$.
  Step 1: Train a deep HSI classifier $\hat{\pi}_{\theta}$ using $\mathcal{D}_{\text{train}}$.
  Step 2: Define the score function $S$ based on $\hat{\pi}_{\theta}$.
  Step 3: Compute the scores for $\mathcal{D}_{\text{cal}}$ and aggregate the results using the Score Aggregation Operator defined in Eq. (5). Denote the aggregated scores as $\{\hat{s}_{i}\}_{i=1}^{{\scriptstyle|\mathcal{D}_{\text{cal}}|}}$.
  Step 4: Determine the threshold $\hat{\tau}$ by:
@@
  Step 5: For $B_{n+1}$, the prediction set is derived as:
@@
To demonstrate that SACP can produce trustworthy prediction sets, we also theoretically establish that SACP ensures coverage, which is formalized in the following proposition.
@@
  • Model-agnostic: SACP is a post-hoc method that can be incorporated into any pre-trained HSI classifier without necessitating modifications to the model itself.
@@
  • Computationally efficient: SACP maintains computational efficiency comparable to that of standard conformal prediction (see the discussion in Section V-C).
@@ V-A Experimental Setup
We utilize the commonly used benchmarks Indian Pines (IP), Pavia University (PU), and Salinas (SA) datasets [^2] These datasets are available at https://www.ehu.eus/ccwintco/index.php/Hyperspectral_Remote_Sensing_Scenes. for our evaluations. Specifically, IP consists of $145\times 145$ pixels and 200 valid spectral bands, representing 16 different land-cover classes. PU includes $610\times 340$ pixels with 103 usable spectral bands, covering 9 land-cover categories. SA comprises $512\times 217$ pixels with 204 effective spectral bands, corresponding to 16 ground-truth classes. For all HSI datasets, we start by randomly sampling HSI cuboids to create the training set. The remaining data is then divided into calibration and test sets according to a pre-defined ratio.
@@
Model $\mathcal{D}_{\mathcal{I}}$ $|\mathcal{D}_{\text{train}}|$ $|\mathcal{D}_{\text{cal}}|$ + $|\mathcal{D}_{\text{test}}|$
1D-CNN IP 250 $9$
PU 103 42673
SA 244 53885
3D-CNN IP 128 10113
PU 52 42724
SA 122 54007
HybridSN IP 128 10113
PU 52 42724
SA 122 54007
SSTN IP 128 10113
PU 52 42724
SA 122 54007
TABLE II: Performance of the pre-trained HSI classifiers used in our experiments.
@@
Pre-trained HSI classifiers
To validate the generalizability of standard conformal prediction and SACP in HSI classification, we employ four widely used HSI classifiers, including both CNN-based and Transformer-based models, each featuring distinct scales and architectures. For CNN-based models, we select 1D-CNN [37], 3D-CNN [38], and HybridSN [39]. The 1D-CNN and 3D-CNN models differ in their convolutional kernel dimensions, while HybridSN leverages multi-dimensional kernels to improve feature extraction. As for the Transformer-based model, we employ SSTN [40]. Each HSI classifier is initially pre-trained on the three datasets and subsequently evaluated on the corresponding calibration and test data. We follow the SSTN training framework available on GitHub [^3] https://github.com/zilongzhong/SSTN for pre-training these classifiers. Table I details the sizes of the training, calibration, and test sets for each dataset across the models. By default, we set the calibration and test sets to be of equal size. Notably, our method remains stable across varying calibration set sizes (see Table IV).
TABLE III: Performance comparison of prediction sets generated by SACP and standard conformal prediction (SCP) across various HSI classifiers with different error rates $\alpha$. Bold numbers indicate that SACP is superior. $\downarrow$ indicates that smaller values are better.
@@
Due to its relatively poor performance, the 1D-CNN is allocated a larger amount of training data compared to the other three models. For clarity, Table II demonstrates the Overall Accuracy (OA) and Average Accuracy (AA) results for the pre-trained HSI classifiers across the three datasets. The definitions of OA and AA are given as follows:
where $\mathcal{D}_{\text{test}}:=\{(B_{i},y_{i})\}_{i=1}^{|\mathcal{D}_{\text{test}}|}$, $\hat{y}_{i}=\mathrm{argmax}(\hat{\pi}_{\theta}(B_{i}))$, $\mathds{1}$ is the indicator function, and $j\in\{1,\dots,|\mathcal{Y}|\}$. Note: In HSI classifier pre-training, the primary focus is not on maximizing their performance metrics such as OA and AA, but rather on achieving model diversity. Our goal is to validate the effectiveness—i.e., the generalizability—of our approach across different model architectures and training processes, irrespective of their performance levels. Given that our primary goal is not to optimize classification performance, we utilize only a minimal amount of data for pre-training these classifiers.
@@
Figure 2: Visualizations of the average size of prediction sets for each pixel in $\mathcal{D}_{\text{cal}}$ and $\mathcal{D}_{\text{test}}$ are presented for the (a) IP and (e) PU datasets. The true labels for each class are shown in (b) and (f). The brightness of each pixel indicates the size of its prediction set, with brighter pixels representing larger sets. Panels (c) and (g) display the set sizes with SACP, while panels (d) and (h) show the set sizes with standard conformal prediction.
@@
V-B Main Results
SACP generates more efficient prediction sets
We compare the performance of prediction sets generated by standard conformal prediction and SACP across various HSI classifiers and datasets with different error rates $\alpha$. The results are presented in Table III, showing that SACP consistently maintains the desired coverage of $1-\alpha$. The results empirically validates the coverage guarantee of SACP, as outlined in Proposition 2. In other words, for each pixel, its true label will be reliably included in the prediction set with the user-specified probabilities. Additionally, SACP consistently produces more efficient prediction sets while maintaining a satisfactory conditional coverage rate. For instance, when applying 3D-CNN to the IP dataset, SACP reduces the size from 5.73 for APS to 3.27, representing a 42.9% decrease. Similarly, for RAPS on the same dataset with 1D-CNN, SACP lowers the size from 3.64 to 2.26, reflecting a 37.9% reduction. These improvements are observed across different error rates $\alpha$, non-conformity score functions, model architectures, and datasets, highlighting the robustness and generalizability of SACP. Notably, unlike traditional classification tasks, SAPS and RAPS do not consistently outperform APS in the context of HSI classification, underscoring the challenges associated with conformal prediction in this field. Nevertheless, SACP consistently improves the performance of various non-conformity score functions, demonstrating its effectiveness.
SACP consistently reduces prediction set size across various pixels
To further demonstrate the effectiveness of SACP, we visualize the size of prediction sets for each pixel of $\mathcal{D}_{\text{cal}}$ and $\mathcal{D}_{\text{test}}$ in Figure 2. This experiment is conducted on the IP and PU datasets using SSTN, with APS as the base non-conformity score function and $\alpha$ set to 0.05. In the visualization, brighter pixels represent larger prediction sets. The results clearly indicate that SACP significantly reduces the size of prediction sets compared to standard conformal prediction. Moreover, standard conformal prediction often produces unusually large prediction sets, even for interior pixels, as evidenced in the lower-left corner of Figure 2(d). Typically, edge pixels, which are more challenging to predict, tend to have larger prediction sets, while interior pixels have smaller sets. SACP effectively addresses this issue by consistently reducing prediction set sizes for both edge and interior pixels. The experimental results further illustrate the superior performance of our approach in conformal HSI classification.